OpenStudy (anonymous):

True or false? If f and g are continuous and f(x) is greater than or equal to g(x) for a≤x≤b then ∫ from a to b f(x)dx ≥ ∫ from a to b g(x) dx. explain your reasoning as to why it's true or false and maybe give an example.

OpenStudy (datanewb):

Well, let's start with the example first. Let's take say: \[g(x) = 3\\f(x) = 7\] Then \[\int_a^bg(x)dx = 3b-3a\] and \[\int_a^bf(x)dx = 7b-7a\], so as long as b > a \[\int_a^bf(x)dx = 7b-7a >3b-3a = \int_a^bg(x)dx\] Also, if you define the integral as the area under the curve from a to b, then it is very clear. (See drawing)